OpenStudy (anonymous):

A. Use order of operations to simplify: B. Compare. 2[3(12-7)____ 5x 5-5 C. Translate the words into a mathematical expression. The number a increased by fifteen is twenty-one. D. Evaluate the expression. xyz – 4z when x = 3, y = 4, and z = 5

OpenStudy (anonymous):

18-3 2-(6-4) = -2592

OpenStudy (phi):

A is missing something it looks like B should have a closing ] like this 2[3(12-7)]

OpenStudy (anonymous):

x*

OpenStudy (anonymous):

That's the first question.

OpenStudy (phi):

18-3 x 2-(6-4) = -2592 ?

OpenStudy (anonymous):

That's what I got unless I;m doing something wrong.

OpenStudy (anonymous):

I'm*

OpenStudy (phi):

how?

OpenStudy (anonymous):

I dont know could you explain all steps

OpenStudy (phi):

did you learn any rules about what to do first? PEDMAS is one way to remember, but there are others

OpenStudy (anonymous):

Yes, I have learned that.

OpenStudy (phi):

P stands for parentheses. Do you have any parentheses?

OpenStudy (anonymous):

Yes

OpenStudy (phi):

can you do the arithmetic inside the parens and replace (6-4) with the answer?

OpenStudy (anonymous):

-24

OpenStudy (phi):

are you saying 6-4 is -24?

OpenStudy (anonymous):

2

OpenStudy (phi):

18-3 x 2-(6-4) now replace (6-4) with your 2 18-3 x 2 - 2 in PEDMAS E is for exponents (I don't see any). D is for divide. nope M is for multiply do you see any numbers to multiply in 18-3 x 2 - 2 ?

OpenStudy (anonymous):

No.

OpenStudy (phi):

doesn't x mean multiply?

OpenStudy (anonymous):

x mines times?

OpenStudy (anonymous):

means

OpenStudy (phi):

yes, or sometimes you see * or just a dot as in \( 3 \cdot 2\)

OpenStudy (anonymous):

So, how would I compare that to B?

OpenStudy (phi):

B is a different question, and does not have anything to do with A. can you finish A? so far you have 18-3 x 2 - 2

OpenStudy (anonymous):

I'm done with A.

OpenStudy (phi):

what did you get?

OpenStudy (anonymous):

18-3 = 15 3x2 = 6 and (6-4) = 2

OpenStudy (phi):

almost. you start with 18-3 x 2-(6-4) you do P (parens) first 6-4 is 2 so replace (6-4) with a 2 to get 18-3 x 2 -2 now do M (multiply) 3x2 is 6 replace 3x2 with 6 to get 18 - 6 -2 now do left to right (18-6) - 2 what do you get?

OpenStudy (anonymous):

-110

OpenStudy (phi):

what is 18-6 ?

OpenStudy (anonymous):

12

OpenStudy (phi):

replace (18-6) with 12 in (18-6) - 2 can you do that?

OpenStudy (phi):

we are almost done. can you rewrite (18-6) - 2 by writing 12 instead of (18-6)

OpenStudy (phi):

yes. 12-2 and finally 10

OpenStudy (phi):

can you make question B in a new post? This one is getting too long